# Functional bottlenecks can emerge from non-epistatic underlying traits

### v1

@@ II.1 Reanalyzing experimental data
In order to precisely answer this question, following Ref. Alqatari and Nagel (2024), we characterize the space of the paths connecting the two wild types upon increasing the threshold $E_{\rm th}$. We consider the subset of paths that (i) are made of single mutations, (ii) connect the red and blue wild types, and (iii) are such that $|E({\bf a})|>E_{\rm th}$ all along the path. We start from $E_{\rm th}=0$, and gradually increase $E_{\rm th}$ to the maximum value such that at least one path remains. The resulting value, which we call $E_{C}$, quantifies how ‘hard’ it is for evolution to find a path that realizes the functionality switch. If $E_{C}\sim E^{WT}$, then one can find a path of single mutations, such that all intermediates are equally functional to the wild types. If instead $E_{C}\ll E^{WT}$, in order to connect the two wild types one has to accept a loss of fitness with respect to the wild types, which makes the transition less likely, but still possible.
@@
For the experimental data of Ref. Poelwijk et al. (2019), this analysis yields a value of approximately $E_{C}\sim 0.53$, or $E_{C}/E^{WT}\sim 0.6$. The resulting space of paths, depicted in Fig. 1d, exhibits the characteristic bottleneck shape also reported in Ref. Alqatari and Nagel (2024), where all paths traverse a single ‘jumper’ genotype with $E=E_{C}$. In Fig. 1d, the jumper genotype has a blue phenotype, after which a single mutation brings to the red phenotype. This functional switch is driven by a mutation located roughly at the midpoint. The number of functional intermediates determines the number of evolutionary paths that can reach the critical ‘jumper’ genotype. More paths imply greater evolutionary accessibility, making it easier for evolution to find the key mutation; while fewer paths indicate a more constrained transition requiring a specific mutational sequence. In contrast, the analysis in Ref. Poelwijk et al. (2019) found a less pronounced bottleneck, with multiple mutational paths surviving - a scenario that can be reproduced here by adopting a less stringent definition of $E_{C}$ which allows for multiple paths, see Ref. Alqatari and Nagel (2024).
@@
The goal of the subsequent analysis is to determine whether this bottleneck structure arises solely from network epistasis or if it can also be explained by a model of global epistasis.
II.2 Toy model of global epistasis
@@
To capture the key ingredients needed for bottleneck topologies, we introduce a minimal toy model featuring only global epistasis, defined as follows:
@@
  • An underlying additive phenotype $E({\bf a})=\sum_{i=1}^{L}h_{i}a_{i}$ has random SMEs $h_{i}$ that are identically and independently distributed according to a symmetric input distribution $P(h)=P(-h)$. After having drawn the values $h_{i}$ independently, we impose that $\sum_{i=1}^{L}h_{i}=0$ by shifting their mean, i.e., $h_{i}\rightarrow h_{i}-\frac{1}{L}\sum_{i=1}^{L}h_{i}$.
@@
  • The two fitness functions that will be associated with the two ‘wild types’ are non-linear functions of the underlying random additive phenotype, $F_{B}=\phi_{B}(E)=\phi_{0}/(1+e^{\beta(E_{\rm th}-E)})$ and $F_{R}=\phi_{R}(E)=\phi_{0}/(1+e^{\beta(E_{\rm th}+E)})$, as illustrated in Fig. 1a. If $E({\bf a})>E_{\rm th}$ then the genotype ${\bf a}$ is functional ‘blue’, if $E({\bf a})<-E_{\rm th}$ it is functional ‘red’, and if $|E({\bf a})|\ll E_{\rm th}$, the genotype is non-functional.
@@
The fitness functions grow sharply around ±$\pm E_{\rm th}$, in a way controlled by the parameter $\beta$. The choice we will make implicitly, following Ref. Alqatari and Nagel (2024), is to consider the limit $\beta\gg 1$, in which the fitness is close to a sharp threshold function at $E_{\rm th}$ (see Fig. 1a, dashed curves). Note that the values $h_{i}$ can be multiplied by an arbitrary constant that can be absorbed into $E_{\rm th}$ and $\beta$, which allows us to fix the overall scale (e.g. the variance) of $P(h)$ without loss of generality.
@@
Given the above definitions, both genotypes ${\bf a}=\mathbf{0}=(0,0,\cdots,0)$ and ${\bf a}=\mathbf{1}=(1,1,\cdots,1)$ have $E=0$ and are thus non-functional. We stress that contrarily to the analysis of experimental data reported above, these two genotypes will not be the wild types in the toy model. Following Ref. Alqatari and Nagel (2024), we thus introduce a ‘tuning procedure’ to generate two wild types, the ‘red’ ${\bf a}_{R}$ with $E_{R}^{WT}<-E_{T}$ and the ‘blue’ ${\bf a}_{B}$ with $E_{B}^{WT}>E_{T}$, for a chosen ‘tuning’ value $E_{T}>0$. The procedure starts from the genotype ${\bf a}=\mathbf{0}=(0,0,\cdots,0)$, thus considered as a ‘common ancestor’, and sequentially introduces mutations as follows.
@@
  • With probability $p$, we perform a ‘greedy’ step. We scan all the $h_{i}$ values that have not been introduced yet (those for which $a_{i}=0$), and choose the one that has the maximum impact towards the desired goal, i.e., the largest positive $h_{i}$ for the blue wild type or the smallest negative $h_{i}$ for the red wild type. We introduce the corresponding mutation by switching $a_{i}=1$.
@@
  • With probability $1-p$, we perform a ‘random’ step. We choose at random one site $a_{i}$ such that $a_{i}=0$, and mutate it to $a_{i}=1$.
The tuning procedure is performed independently for the two wild types starting from ${\bf a}=\mathbf{0}$, and continues until for the first time $E^{WT}_{R}<-E_{T}$ to construct the red wild type, or $E^{WT}_{B}>E_{T}$ to construct the blue wild type. The motivation for this procedure is to mimic somehow the tuning of elastic networks performed in Ref. Alqatari and Nagel (2024), in which the network is tuned based on a local proxy for the global fitness; we hereby assume that this local proxy can indeed provide the best choice (with probability $p$) or miss it (with probability $1-p$).
The parameters that define the model are the sequence length $L$, the a priori distribution of SMEs $P(h)$, the probability $p$ and the tuning parameter $E_{T}$ that controls the wild type construction. We found that many choices of $P(h)$ lead to similar outcomes if the model is properly calibrated. For simplicity, we thus restrict to two natural choices.
@@
[…]
For a given choice of parameters, an ‘instance’ of our random model is thus defined by the set of the $L$ SMEs, $h_{i}$, and by the two wild types ${\bf a}_{R}$ and ${\bf a}_{B}$. For a given instance, we can define the number of mutations $M$ separating ${\bf a}_{R}$ and ${\bf a}_{B}$ (i.e., the number of $a_{i}$ that differ in the two sequences), and we can then investigate the $M!$ directed paths that connect them, obtained by introducing these mutations in all possible orders. Following Ref. Alqatari and Nagel (2024) and as discussed above, we can find the maximum value of $E_{\rm th}$ such that at least one single-mutational path connecting the two wild types, with all intermediate states having $|E|>E_{\rm th}$, survives, and we call this value $E_{C}$. The closer $E_{C}$ is to $E^{WT}$, the easier it is to perform the functional switch.
Given $(L,p,E_{T})$ and $P(h)$, we wish to characterize the statistical properties of the quantities $M$, $E^{WT}$, and $E_{C}$, which characterize the overall topology of the fitness landscape. We will also define $E^{WT}_{\rm max}=\max(E^{WT}_{B},|E^{WT}_{R}|)$. In the following, we will use brackets, $\langle A\rangle$, to denote the statistical average of an observable $A$ over an ensemble of random instances generated by the model. Unless otherwise specified, the average is taken over 20000 independent instances of the model for each value of the parameters.
Our model is similar in spirit to Fisher’s geometric model (FGM) as discussed in Refs. De Visser and Krug (2014); Hwang et al. (2017); Pahujani and Krug (2025), from which it however differs in two important aspects. First, the choice of non-linearity, which reflects here the existence of two distinct fitness functions associated to each wild type. Second, the fact that we choose two reference wild types using the stochastic procedure described above, and we then restrict to the space of intermediates between the selected wild types. As we show in the following, these two ingredients, together with a proper calibration of the model, give rise to the desired bottleneck structure.
Before discussing the details of the calibration procedure, a typical example of the space of mutational paths obtained from a calibrated model is shown in Fig. 2. The similarity with Fig. 1d, modulo the (irrelevant) change of scale of $E$, is striking.
@@ II.3 Calibration of the model
Figure 2: Same representation as in Fig. 1d, using an instance of the calibrated model with Gaussian $P(h)$ instead of the experimental data. The lower panel reports the value of $E$ for each functional variant as a function of the distance from the blue wild type (here with $E^{WT}_{B}/E_{T}\approx 1.50$ and $E^{WT}_{R}/E_{T}\approx-1.79$). The threshold value for which at least a mutational path remains is $E_{C}/E_{T}\approx 0.72$ (dashed lines) and when normalized with respect to the largest of the two wild-types it reads $E_{C}/E^{WT}_{\max}\approx 0.40$.
@@
Figure 3: Calibration of the two models with different $P(h)$, a Gaussian distribution or a Pareto distribution with cutoff. The black symbols correspond to the choice of $p$ after calibration. (a) Value of $E_{T}$ for which $\langle M\rangle\approx 8$ as a function of $p$. Dashed lines are linear fits. (b) Average value $\langle E^{WT}_{B}/E_{T}\rangle$ as a function of $p$. (c) Average of $E_{C}/E^{WT}_{\max}$ as a function of $p$. (d) Probability of having a single jumper at $E_{C}$, or equivalently that $E_{C}<E^{WT}_{\min}$, as a function of $p$.
@@
The first step is to calibrate the parameters $(L,p,E_{T})$ in such a way that our toy model reproduces the basic phenomenology observed in Refs. Poelwijk et al. (2019); Alqatari and Nagel (2024). The requirements are the following.
@@
  • The phenotype of the two wild types should be comparable to $E_{T}$, i.e. $|E^{WT}|/E_{T}\simeq 1$, in order for $E_{T}$ to be a meaningful fitness scale.
@@
  • We want to reproduce a ‘bottlenecked’ structure of the space of evolutionary paths connecting the two wild types, as shown in Figs. 1d, 2 and in Ref. Alqatari and Nagel (2024). This is characterized by a single ‘jumper’ genotype through which all paths must pass. The mutation responsible for the functionality switch occurs just after (or just before) this genotype. We note that the jumper genotype can be connected by a single mutation to only one, or more than one, genotypes carrying the other phenotype.
@@
  • The ‘jumper’ genotype should occur approximately at half distance along the evolutionary trajectory, as it is observed in experimental data (Fig. 1d). Furthermore, we want the largest possible number of functional intermediate variants to survive before and after the jump, thus maximizing the number of functional evolutionary paths connecting the two wild types.
@@
  • The fitness variation produced by the ‘jumper’ mutation(s) should be as large as possible, in such a way that $E_{C}$ is as close as possible to $E^{WT}$, as discussed in Ref. Alqatari and Nagel (2024). This makes the resulting mutational paths more viable for evolution.
@@
  • Finally, for practical reasons, the number of mutations $M$ connecting the two wild types should be small enough, otherwise enumerating all the $M!$ paths becomes computationally very expensive. To avoid this, we impose $\langle M\rangle\sim 8$, a value close to that of Ref. Alqatari and Nagel (2024).
@@
An important remark is in order about the expected value of $E_{C}$. Consider for illustration the case in which the tuning procedure reaches the target with a single greedy step for both wild types. We can thus expect $E^{WT}_{B}\sim\max(h_{i})$ and $E^{WT}_{R}\sim\min(h_{i})$. In absence of network epistasis, the largest jumper mutation that can happen along a path has $|\Delta E|\sim\max(|h_{i}|)\sim E^{WT}_{\rm max}$, but this mutation must bring the system from $E_{C}$ to $-E_{C}$, hence $2E_{C}\sim|\Delta E|\sim E^{WT}_{\rm max}$. Based on this simple argument, the largest possible value of $E_{C}$ we can expect is such that $E_{C}/E^{WT}_{\rm max}\sim 1/2$, as we observe numerically (Fig. 3c).
@@
To fix the model parameters $(L,p,E_{T})$ we proceed as follows. First of all, we fix $L=500$ as the length of a typical protein. We checked that the results are very weakly dependent on $L$ (see the SI). Next, for each $p$, we progressively increase $E_{T}$ until the average number of mutations $\langle M\rangle$ separating the two wild types is $\langle M\rangle\sim 8$. The resulting $E_{T}(p)$ is plotted as a function of $p$ for both choices of $P(h)$ in Fig. 3a, and we find that it is approximately linear in $p$. The linear fit allows us to fix $E_{T}(p)$ for all $p$.
@@
The next step is to calibrate $p$. For this, we consider the ratio $|E^{WT}|/E_{T}$ for both wild types, which we want to be close to one. The evolution of its statistical mean with $p$ is shown in Fig. 3b. For small $p$, we mostly perform random steps in the training, and the requirement that $M$ is small forces $E_{T}$ to be small as well (Fig. 3a). As a result, it is very likely to ‘overshoot’ the target and end up with $|E^{WT}|/E_{T}\gg 1$. Hence, in order to have $|E^{WT}|/E_{T}\sim 1$, we need a large enough $p$.
@@
Next, we consider the connectivity threshold $E_{C}$. We recall that $E_{C}$ is obtained, for each instance, in such a way that at least one single-mutational path connecting the two wild types exists, with all intermediate states having $|E|>E_{C}$. Because both wild types must always be included in the space of allowed genotypes, increasing the value of $E_{C}$ beyond $E^{WT}_{\min}=\min(E^{WT}_{B},|E^{WT}_{R}|)$ is not meaningful Alqatari and Nagel (2024). This gives an upper bound for the largest value of $E_{C}$ for each instance. We show in Fig. 3c its average value as a function of $p$, where the average is restricted to instances such that $E_{C}<0.9E^{WT}_{\min}$. When $p$ is too large, the tuning procedure requires too many greedy steps and as a consequence $\langle E_{C}/E^{WT}_{\rm max}\rangle$ tends to decrease, contrarily to what we want in order to have bottleneck topologies with large $E_{C}$. This suggests to lower the value of $p$ as much as possible.
@@
However, in Fig. 3d we report the probability of having a single jumper at $E_{C}$, which we estimate for numerical convenience by the probability that $E_{C}<0.9E^{WT}_{\min}$, as a function of $p$. When $p$ is too small, this probability drops quickly below one. This is because $|E^{WT}|/E_{T}$ starts increasing, the wild types are obtained by many random steps, and the bottleneck structure is lost.
@@
Taken together, these results suggest to choose the lowest value of $p$ such that the probability of a single jumper is still very close to one, which also gives both $E^{WT}/E_{T}\sim 1$ and $E_{C}/E^{WT}_{\rm max}\sim 0.5$. The value of $E_{T}(p)$ is fixed by the requirement that $\langle M\rangle\sim 8$. For both choices of $P(h)$, this optimal value is around $p=0.25$, although slightly larger values could be considered as well.
@@
The crucial insight from the calibration is that one needs to tune $p$ in order to achieve a good balance between greedy and random steps. Correspondingly, as we will discuss later, the distribution of the SMEs that are selected by the tuning procedure that generates the wild types is very heterogeneous, with a coexistence of large SMEs (generated by greedy steps) and small SMEs (generated by random steps).

### v2

@@ II.1 Experimental data on fitness bottlenecks
In order to precisely answer this question, following Ref. Alqatari and Nagel (2026), we characterize the space of the paths connecting the two reference variants upon increasing the threshold $E_{\rm th}$. We consider the subset of paths that (i) are made of single mutations, (ii) connect the red and blue reference sequences, and (iii) are such that $|E({\bf a})|>E_{\rm th}$ all along the path. We start from $E_{\rm th}=0$, and gradually increase $E_{\rm th}$ to the maximum value such that at least one path remains. The resulting value, which we call $E_{C}$, quantifies how ‘hard’ it is for evolution to find a path that realizes the functionality switch. If $E_{C}\sim E^{\rm{ref}}$, then one can find a path of single mutations, such that all intermediates are equally functional to the reference variants. If instead $E_{C}\ll E^{\rm{ref}}$, in order to connect the two reference variants one has to accept a loss of fitness with respect to the reference fitness, which makes the transition less likely, but still possible.
@@
For the experimental data of Ref. Poelwijk et al. (2019), this analysis yields a value of approximately $E_{C}\sim 0.53$, or $E_{C}/E^{\rm{ref}}\sim 0.6$. The resulting space of paths, depicted in Fig. 1d, exhibits the characteristic bottleneck shape also reported in Ref. Alqatari and Nagel (2026), where all paths traverse a single ‘jumper’ genotype with $E=E_{C}$. In Fig. 1d, the jumper genotype has a blue phenotype, after which a single mutation brings to the red phenotype. This functional switch is driven by a mutation located roughly at the midpoint. The number of functional intermediates determines the number of evolutionary paths that can reach the critical ‘jumper’ genotype. More paths imply greater evolutionary accessibility, making it easier for evolution to find the key mutation; while fewer paths indicate a more constrained transition requiring a specific mutational sequence. Note that the analysis in Ref. Poelwijk et al. (2019) found a less pronounced bottleneck, with multiple mutational paths surviving – a scenario that can be reproduced here by adopting a less stringent definition of $E_{C}$ which allows for multiple paths, see Ref. Alqatari and Nagel (2026).
II.2 Stylized model of global epistasis
@@
The goal of this paper is to determine whether a bottleneck structure like the one observed in Sec. II.1 requires network epistasis, or if it can also be explained by a simpler model of global epistasis.
To formulate a simple stylized model, we begin by an important observation. The reference proteins analyzed in Ref. Poelwijk et al. (2019) are not the product of natural evolution, but rather highly engineered sequences. In engineered proteins, derived either by random mutagenesis as in Sec. II.1 or by directed evolution, one typically observes that the fitness improvement is achieved by just a few highly beneficial mutations, see e.g. Zhong et al. (2020); Buda et al. (2023); Fröhlich et al. (2024). While this might seem, in principle, a highly specific feature of engineered proteins, a similar scenario has been observed when a natural protein needs to quickly adapt to a novel environment and acquire a new function. A notable example is the SARS-COV-2 Spike protein that displayed a few highly beneficial mutations just after the virus jumped to human hosts (see e.g. Rodriguez-Rivas et al. (2022)), but similar dynamics has been observed in other viral proteins Hadfield et al. (2018), and in the immune system Phillips et al. (2021). In each of these cases, just a few advantageous mutations allow the protein to acquire the new desired function, whether the selection pressure is applied artificially or emerges from a change in the natural environment. A very different dynamics might be at play in the neutral evolution of optimized initial proteins Fantini et al. (2020); Stiffler et al. (2020); Bisardi et al. (2022); Bari et al. (2024), which might explain why the combinatorial landscape of Ref. Westmann et al. (2024) does not display a bottleneck.
@@
To capture a few minimal ingredients inspired by the results of Sec. II.1 and the above discussion, we introduce a stylized model featuring only global epistasis. The model is defined as follows:
@@
  • An underlying additive trait $E({\bf a})=\sum_{i=1}^{L}h_{i}a_{i}$ has random SMEs $h_{i}$ that are identically and independently distributed according to a symmetric input distribution P(h)=P(h)P(h)=P(-h). After having drawn the values $h_{i}$ independently, we impose that $\sum_{i=1}^{L}h_{i}=0$ by shifting their mean, i.e., $h_{i}\rightarrow h_{i}-\frac{1}{L}\sum_{i=1}^{L}h_{i}$.
@@
  • The two fitness functions that will be associated with the two ‘colors’ are non-linear functions of the underlying trait, $F_{B}=\phi_{B}(E)=\phi_{0}/(1+e^{\beta(E_{\rm th}-E)})$ and $F_{R}=\phi_{R}(E)=\phi_{0}/(1+e^{\beta(E_{\rm th}+E)})$, as illustrated in Fig. 1a. If $E({\bf a})>E_{\rm th}$ then the genotype $\bf a$} is functional ‘blue’, if $E({\bf a})<-E_{\rm th}$ it is functional ‘red’, and if $|E({\bf a})|\ll E_{\rm th}$, the genotype is non-functional.
@@
The fitness functions grow sharply around $\pm E_{\rm th}$, in a way controlled by the parameter $\beta$. The choice we will make implicitly, following Ref. Alqatari and Nagel (2026), is to consider the limit $\beta\gg 1$, in which the fitness is close to a threshold (or Heaviside) function at $\pm E_{\rm th}$ (see Fig. 1a, dashed curves). Note that the values $h_{i}$ can be multiplied by an arbitrary constant that can be absorbed into $E_{\rm th}$ and $\beta$, which allows us to fix the overall scale (e.g. the variance) of $P(h)$ without loss of generality.
@@
Given the above definitions, both genotypes ${\bf a}=\mathbf{0}=(0,0,\cdots,0)$ and ${\bf a}=\mathbf{1}=(1,1,\cdots,1)$ have $E=0$ and are thus non-functional. We stress that, contrarily to the analysis of experimental data reported above, these two genotypes will not be the reference sequences in the stylized model. Following Ref. Alqatari and Nagel (2026), we thus introduce a ‘tuning procedure’ to generate two reference variants, the ‘red’ ${\bf a}_{R}$ with $E_{R}^{\rm{ref}}<-E_{T}$ and the ‘blue’ ${\bf a}_{B}$ with $E_{B}^{\rm{ref}}>E_{T}$, for a chosen ‘tuning’ value $E_{T}>0$.
The procedure starts from the genotype ${\bf a}=\mathbf{0}=(0,0,\cdots,0)$, which is thus considered as a ‘common ancestor’ or ‘wild type’ (for example, in the experiment described in Sec. II.1, it would correspond to the eqFP578 protein). One then sequentially introduces mutations as follows.
@@
  • With probability $p$, we perform a ‘greedy’ step. We scan all the $h_{i}$ values that have not been introduced yet (those for which $a_{i}=0$), and choose the one that has the maximum impact towards the desired goal, i.e., the largest positive $h_{i}$ for the blue reference variant or the smallest negative $h_{i}$ for the red reference variant. We introduce the corresponding mutation by switching $a_{i}=1$.
@@
  • With probability 1p1-p, we perform a ‘random’ step. We choose at random one site $a_{i}$ such that $a_{i}=0$, and mutate it to $a_{i}=1$.
The tuning procedure is performed independently for the two colors starting from ${\bf a}=\mathbf{0}$, and continues until for the first time $E^{\rm{ref}}_{R}<-E_{T}$ to construct the red reference variant, or $E^{\rm{ref}}_{B}>E_{T}$ to construct the blue reference variant.
This procedure is not designed to be a fully realistic model of molecular evolution. Nonetheless, it is grounded in experimental observations. As noted at the beginning of this section, directed evolution experiments often yield final protein variants carrying a small number of beneficial mutations together with a few neutral ones. More generally, when a protein is challenged to acquire a new function, adaptation often involves selecting a handful of advantageous substitutions on top of incidental neutral changes. Our approach for constructing the reference variants is simply the most straightforward way to reproduce this pattern, without any claim of population-genetic realism. We also believe that this procedure mimics the tuning of elastic networks performed in Ref. Alqatari and Nagel (2026), in which the network is tuned based on a local proxy for the global fitness; we hereby assume that this local proxy can indeed provide the best choice (with probability $p$) or miss it (with probability 1p1-p).
The parameters that define the model are the sequence length $L$, the a priori distribution of SMEs $P(h)$, the probability $p$ and the tuning parameter $E_{T}$ that controls the reference sequence construction. We note that $P(h)$ may be interpreted as describing the a priori distribution of mutational effects – arising, for instance, from underlying biophysical constraints – which represents the potential ’pool’ of mutations accessible to evolution. We found that various choices of $P(h)$ lead to similar outcomes, provided the model is appropriately calibrated. For the sake of parsimony, and recognizing that these are stylized representations, we restrict our analysis to two distributions that capture distinct mutational behaviors.
@@
    Pareto cutoff: Motivated by the analysis of experimental data in Fig. 1c, a Pareto distribution, i.e., Eq. (6) with $x=h$, where we fix $\alpha=0.7$ and $x_{m}=0.1$. We also include an upper cutoff as in the data, i.e. we set $P(h)=0$ for $|h|>2$ and adjust the normalization constant accordingly.
For a given choice of parameters, an ‘instance’ of our random model is thus defined by the a priori pool of the $L$ SMEs, $h_{i}$, and by the two reference variants ${\bf a}_{R}$ and ${\bf a}_{B}$. For a given instance, we can define the number of mutations $M$ separating ${\bf a}_{R}$ and ${\bf a}_{B}$ (i.e., the number of $a_{i}$ that differ in the two sequences), and we can then investigate the $M!$ directed paths that connect them, obtained by introducing these mutations in all possible orders. Following Ref. Alqatari and Nagel (2026) and as discussed above, we can find the maximum value of $E_{\rm th}$ such that at least one single-mutational path connecting the two reference variants, with all intermediate states having $|E|>E_{\rm th}$, survives, and we call this value $E_{C}$. The closer $E_{C}$ is to $E^{\rm{ref}}$, the easier it is to perform the functional switch.
Given $(L,p,E_{T})$ and $P(h)$, we wish to characterize the statistical properties of the quantities $M$, $E^{\rm{ref}}$, and $E_{C}$, which characterize the overall topology of the fitness landscape. Another interesting quantity is the distribution P(h~)P($\tilde${h}) of SMEs $\tilde{h}_{i}$ selected by the tuning procedure to generate the two reference variants (i.e., those $h_{i}$ that correspond to $a_{i}=1$ in each reference genotype), which may be interpreted as the distribution of mutations that are ‘fixed’ by evolution. We will also define $E^{\rm{ref}}_{\rm max}=\max(E^{\rm{ref}}_{B},|E^{\rm{ref}}_{R}|)$. In the following, we will use brackets, $\langle A\rangle$, to denote the statistical average of an observable $A$ over an ensemble of random instances generated by the model. Unless otherwise specified, the average is taken over 20000 independent instances of the model for each value of the parameters.
Our model is similar in spirit to Fisher’s geometric model (FGM) as discussed in Refs. De Visser and Krug (2014); Tenaillon (2014); Martin and Lenormand (2015); Hwang et al. (2017); Pahujani and Krug (2025), from which it however differs in two important aspects. First, the choice of non-linearity, which reflects here the existence of two distinct fitness functions associated to each color, see also Ref. Martin and Lenormand (2015) for a similar choice. Second, the fact that we choose two reference variants using the stochastic procedure described above, and we then restrict to the space of intermediates between the selected variants. As we show in the following, these two ingredients, together with a proper calibration of the model, give rise to the desired bottleneck structure.
Before discussing the details of the calibration procedure, a typical example of the space of mutational paths obtained from a calibrated model is shown in Fig. 2. The qualitative similarity with Fig. 1d, modulo the (irrelevant) change of scale of $E$, is striking.
@@ II.3 Calibration of the model
Figure 2: Same representation as in Fig. 1d, using an instance of the calibrated model with Gaussian $P(h)$ instead of the experimental data. The lower panel reports the value of $E/E_{T}$ for each functional variant as a function of the distance from the blue reference variant (here with $E^{\rm{ref}}_{B}/E_{T}\approx 1.50$ and $E^{\rm{ref}}_{R}/E_{T}\approx-1.79$). The threshold value for which at least a mutational path remains is $E_{C}/E_{T}\approx 0.72$ (dashed lines) and when normalized with respect to the largest of the two reference genotypes it reads $E_{C}/E^{\rm{ref}}_{\max}\approx 0.40$.
@@
Figure 3: Calibration of the two models with different $P(h)$, a Gaussian distribution or a Pareto distribution with cutoff. The black symbols correspond to the choice of $p$ after calibration. (A) Value of $E_{T}$ for which $\langle M\rangle\approx 8$ as a function of $p$. Dashed lines are linear fits. (B) Average value $\langle E^{\rm{ref}}_{B}/E_{T}\rangle$ as a function of $p$. (C) Average of $E_{C}/E^{\rm{ref}}_{\max}$ as a function of $p$. (D) Probability of having a single jumper at $E_{C}$, or equivalently that $E_{C}<E^{\rm{ref}}_{\min}$, as a function of $p$.
@@
The first step is to calibrate the parameters $(L,p,E_{T})$ in such a way that our stylized model reproduces the basic phenomenology observed in Refs. Poelwijk et al. (2019); Alqatari and Nagel (2026). The requirements are the following.
@@
  • The phenotype of the two reference sequences should be comparable to $E_{T}$, i.e. $|E^{\rm{ref}}|/E_{T}\simeq 1$, in order for $E_{T}$ to be a meaningful fitness scale.
@@
  • We want to reproduce a ‘bottlenecked’ structure of the space of evolutionary paths connecting the two reference variants, as shown in Figs. 1d, 2 and in Ref. Alqatari and Nagel (2026). This is characterized by a single ‘jumper’ genotype through which all paths must pass. The mutation responsible for the functionality switch occurs just after (or just before) this genotype. We note that the jumper genotype can be connected by a single mutation to only one, or more than one, genotypes carrying the other phenotype.
@@
  • The ‘jumper’ genotype should occur approximately at half distance along the evolutionary trajectory, as it is observed in experimental data (Fig. 1d). Furthermore, we want the largest possible number of functional intermediate variants to survive before and after the jump, thus maximizing the number of functional evolutionary paths connecting the two reference variants.
@@
  • The fitness variation produced by the ‘jumper’ mutation(s) should be as large as possible, in such a way that $E_{C}$ is as close as possible to $E^{\rm{ref}}$. If this is the case, a path can go in a single jump from having fitness close to the red reference, to having fitness close to the blue reference. Otherwise, one would necessarily have to tolerate a decrease in fitness to perform the functional switch. This makes the resulting mutational paths more viable for evolution.
@@
  • Finally, for practical reasons, the number of mutations $M$ connecting the two reference variants should be small enough, otherwise enumerating all the $M!$ paths becomes computationally very expensive. To avoid this, we impose $\langle M\rangle\sim 8$, a value close to that chosen in Ref. Alqatari and Nagel (2026).
@@
An important remark is in order about the expected value of $E_{C}$. Consider for illustration the case in which the tuning procedure reaches the target with a single greedy step for both colors. We can thus expect $E^{\rm{ref}}_{B}\sim\max(h_{i})$ and $E^{\rm{ref}}_{R}\sim\min(h_{i})$. In absence of network epistasis, the largest jumper mutation that can happen along a path has $|\Delta E|\sim\max(|h_{i}|)\sim E^{\rm{ref}}_{\rm max}$, but this mutation must bring the system from $E_{C}$ to $E_{C}$, hence $2E_{C}\sim|\Delta E|\sim E^{\rm{ref}}_{\rm max}$. Based on this simple argument, the largest possible value of $E_{C}$ we can expect is such that $E_{C}/E^{\rm{ref}}_{\rm max}\sim 1/2$, as we observe numerically (Fig. 3c).
@@
To fix the model parameters $(L,p,E_{T})$ we proceed as follows. First of all, we fix $L=500$ as the length of a typical protein. We checked that the results are very weakly dependent on $L$ (see the SI). Next, for each $p$, we progressively increase $E_{T}$ until the average number of mutations $\langle M\rangle$ separating the two reference variants is $\langle M\rangle\sim 8$. The resulting $E_{T}(p)$ is plotted as a function of $p$ for both choices of $P(h)$ in Fig. 3a, and we find that it is approximately linear in $p$. The linear fit allows us to fix $E_{T}(p)$ for all $p$.
@@
The next step is to calibrate $p$. For this, we consider the ratio $|E^{\rm{ref}}|/E_{T}$ for both colors, which we want to be close to one. The evolution of its statistical mean with $p$ is shown in Fig. 3b. For small $p$, we mostly perform random steps in the training, and the requirement that $M$ is small forces $E_{T}$ to be small as well (Fig. 3a). As a result, it is very likely to ‘overshoot’ the target and end up with $|E^{\rm{ref}}|/E_{T}\gg 1$. Hence, in order to have $|E^{\rm{ref}}|/E_{T}\sim 1$, we need a large enough $p$.
@@
Next, we consider the connectivity threshold $E_{C}$. We recall that $E_{C}$ is obtained, for each instance, in such a way that at least one single-mutational path connecting the two reference variants exists, with all intermediate states having $|E|>E_{C}$. Because both reference sequences must always be included in the space of allowed genotypes, increasing the value of $E_{C}$ beyond $E^{\rm{ref}}_{\min}=\min(E^{\rm{ref}}_{B},|E^{\rm{ref}}_{R}|)$ is not meaningful Alqatari and Nagel (2026). This gives an upper bound for the largest value of $E_{C}$ for each instance. We show in Fig. 3c its average value as a function of $p$, where the average is restricted to instances such that $E_{C}<0.9E^{\rm{ref}}_{\min}$. When $p$ is too large, the tuning procedure requires too many greedy steps and as a consequence $\langle E_{C}/E^{\rm{ref}}_{\rm max}\rangle$ tends to decrease, contrarily to what we want in order to have bottleneck topologies with large $E_{C}$. This suggests to lower the value of $p$ as much as possible.
@@
However, in Fig. 3d we report the probability of having a single jumper at $E_{C}$, which we estimate for numerical convenience by the probability that $E_{C}<0.9E^{\rm{ref}}_{\min}$, as a function of $p$. When $p$ is too small, this probability drops quickly below one. This is because $|E^{\rm{ref}}|/E_{T}$ starts increasing, the reference variants are obtained by many random steps, and the bottleneck structure is lost.
@@
Taken together, these results suggest to choose the lowest value of $p$ such that the probability of a single jumper is still very close to one, which also gives both $E^{\rm{ref}}/E_{T}\sim 1$ and $E_{C}/E^{\rm{ref}}_{\rm max}\sim 0.5$. The value of $E_{T}(p)$ is fixed by the requirement that $\langle M\rangle\sim 8$. For both choices of $P(h)$, this optimal value is around $p=0.25$, although slightly larger values could be considered as well.
@@
The crucial insight from the calibration is that one needs to tune $p$ in order to achieve a good balance between greedy and random steps. Correspondingly, as we will discuss later, the distribution P(h~)P($\tilde${h}) of the SMEs that are selected by the tuning procedure that generates the reference sequences (that may be interpreted as those ‘fixed by evolution’) is very heterogeneous, featuring a coexistence of large SMEs (generated by greedy steps) and small SMEs (generated by random steps).